OpenStudy (anonymous):

Find the length of the hypotenuse of an isosceles right triangle whose legs are 1 unit in length

OpenStudy (btaylor):

\[1^2 + 1^2 = c^2 \rightarrow c = \sqrt{2}\]

OpenStudy (anonymous):

is that a \[\sqrt{^2} or \sqrt{2}\]

OpenStudy (btaylor):

it is the square root of 2.